Find an equation of the line. Write the equation using function notation.
(2,-6)
​;
parallel to 5x+2y= 7
  f(x)=

Answer by stanbon(75887)

Find an equation of the line. Write the equation using function notation.
(2,-6); parallel to 5x+2y= 7
Find the slope of the given line.
5x + 2y = 7
2y = -5x+7
y = (-5/2)x + (7/2)
slope = -5/2
------
Form of the answer:: y = mx + b
Solve for "b"::
-6 = (-5/2)2 + b
-6 = -5 + b
b = -1
---
Ans: y = (-5/2)x - 1
----------------
Cheers,
Stan H.
------------